Given that; P= 144-12Q

i) Determine the equation for the Total revenue and Marginal revenue

ii) At what output is Marginal Revenue = 0

iii) At what output is Total revenue at maximum

iv) Determine the price elasticity of demand at output where Total Revenue is maximum



1
Expert's answer
2021-08-10T11:43:10-0400
"TR=pQ=144Q-12Q^2"

"MR=144-24Q"

"144-24Q=0"

"Q=6"

"p=144-12\times6=72"

"E=-12\times \frac{72}{6}=-144"